J. McWhorter of College of Health Science at the University of Nevada studied physical therapy students during their graduate-school year. The researchers were interested in the fact that, although graduate physical therapy students are taught the principles of fitness, some have difficulty finding time to implement those principles. A sample of 27 female graduate physical-therapy students had a mean of 22.46 percent body fat.
Assuming the percent body fat of females graduate physical-therapy students is normally distributed with a standard deviation of 4.10 percent body fat, determine a 95% confidence interval for the mean percent body fat of all female graduate physical-therapy students.

Answer by stanbon(75887)

A sample of 27 female graduate physical-therapy students had a mean of 22.46 percent body fat.
Assuming the percent body fat of females graduate physical-therapy students is normally distributed with a standard deviation of 4.10 percent body fat, determine a 95% confidence interval for the mean percent body fat of all female graduate physical-therapy students.
------
x-bar = 22.46
ME = z*s/sqrt(n) = 1.96*4.1/sqrt(27) = 1.547
------
95%CI:: 22.46-1.547 < u < 22.46+1.547
---------------------
Cheers,
Stan H.
--------------